March 5, 2023

From HallCWiki
Jump to: navigation, search

Return to the calendar.

RC Meeting Information

Current RC: Iuliia Skorodumina.

No RC meeting today.


Saturday Day (03/04). The shift started with the Accelerator issues. 2L26-5 ioc went down ( At around 12:30 the issue was fixed, but in about two hours, the same ioc went down again ( No beam for the rest of the shift. Day Shift summary: .

Saturday Swing (03/04). The beam resumed about 7pm and since then was quite stable. We were taking production data at 60 uA. Swing Shift summary: .

Sunday Owl (03/05). We were taking data overnight at 60 uA. A quiet and productive shift with large ABU: .

For the current setting (Pm=800 MeV/c), we have collected a total of 9655 mC integrated charge out of the planned 15696 mC, which corresponds to 62%.

Accelerator Status

The issue with Hall A BLM trips still persists. From Friday's studies, the conclusion was made that the 4th pass separator is likely responsible for that ( Unfortunately, this is not an easy fix and the repair work will realistically take a day and a half. The repair is scheduled for next Tuesday (March 7). For the weekend, it was decided that Halls B, C, and D will be getting beam, while Hall A will wait until Tuesday's repair.

Current Run Plan

The plan is to continue taking production data for the Pm=800 MeV/c at the highest possible current.